Find D.
A political pollster is conducting an analysis of sample results in order to make predictions on election night. Assuming a two-candidate election, if a specific candidate receives at least 55% of the vote in the sample, that candidate will be forecast as te winner of he election. You select a random sample of 100 votes. Complete parts (a) through (c) below Home ork The probability is .1151 that a candidate wil be forecast as the winner when the population percentage of her vote is 49% & Tes Round to four decimal places as noeded.) d. Suppose that the sample size was increased to 400. Repeat process (a) through (c), using this new sample size. Comment on the difference. en The probability is .0250 that a candidate wil be forecast as the winner when the populason percentage of her vote is 50.1% (Round to four decimal places as needed.) ook The probability is that a candidate wil be forecast as the winner when the population percentage of her vote is 58% r Cont (Round to four decimal places as needed.) or Suce Enter your answer in the answer box and then click Check Answer nedia L remaining Clear All Check Answer eTools
Answer #1

d) Calculate the probability that the candidate wil be forecast as the winner when the population percentage of her vote is 50.1% p-p p(1-p 72 , 0.55-p p(1-p P(B2055)-1- 72 0.55-0.501 0.501(1-0.501) 400 1-P(Z<1.96) 1-0. 9750 0.0250 The probability is 0.0250 that a candidate will be forecast as the winner when the population percentage of her vote is 50.1% Calculate the probability that the candidate wil be forecast as the winner when the population percentage of her vote is 58%P(p2058) = 1-P- p-p p(1-p p(1-p 72 72 1-P Z <-0.55-0.58 0.58(1-0.58) 400) = 1-P(Z <-1.22) -1-0.1121 -0.8879 The probability is 0.8879 that a candidate will be forecast as the winner when the population percentage of her vote is 58%